OpenStudy (anonymous):

Solve: negative (1 over 2 x ) + 3 = −x + 7

OpenStudy (anonymous):

-1/2x +3 = -x +7 Multiply throughout by 2x -1 + 6x = -2x^2 + 14x 2x^2 - 8x -1 =0 Factorize

OpenStudy (anonymous):

how did you get negative? for 1/2